A bowling ball with a mass of 15.0 kg is rolling down a hill that is inclined at an angle of 16 degrees. The coefficient of kinetic friction between the ball and the grass is 0.10. Calculate the velocity of the bowling ball obtained after 8 seconds and the distance travelled during that 8 seconds.

1
Expert's answer
2022-03-04T08:31:13-0500

"a=g(\sin\alpha-\mu\cos\alpha)=0.18~\frac m{s^2},"

"v=at=1.44~\frac ms,"

"s=\frac{at^2}2=5.7~m."